A car starts from rest. It acquires a speed of 25 metre per second after 20 seconds find distance travelled by the car.


1
Expert's answer
2021-02-10T21:37:25-0500
"d=\dfrac{v_0+v}{2}t,""d=\dfrac{0+25\ \dfrac{m}{s}}{2}\cdot20\ s=250\ m."